Find
∫e^1/3dx

3e^x/3+c
ex+c
3e1/3+c
1/3ex+c
1
Expert's answer
2014-09-12T14:22:47-0400
∫e^1/3dx=e^1/3∫dx=(e^1/3)*x+C, where C is an arbitrary real constant.